NEED ASAP PLEASE WILL GIVE GOOD RATE RIGHT AWAY FOR ALL ANSWERS!
in openGL
given a pixel represented as a triple float color values (0.3,0.5,0.8) convert the pixel to byte color values . Note the range of a byte color value is [0,255] and the range of a float color is [0.0,1.0]. Round answer to the nearest whole number
Answer #1

The formula to convert an float color value to byte color value is as follows:

b(n) = f(n) * 255 ; where b(n) is the byte color value of 'n' and f(n) is the float color value of 'n', where n is an real number

Now the byte value for 0.3 is

b(0.3) = 0.3 * 255 = 76.5

b(05) = 0.5 * 255 = 127.5

b(0.8) = 0.8 * 255 = 204

Now the nearest whole number to 76.5 is 77

And the nearest whole number to 127.5 is 128

And the nearest whole number to 204 is 204

So the float color values (0.3,0.5,0.8) in byte color values to the nearest whole number are (77,128,204)